A Holistic View of Lexical Choice

Stephan Busemann^{∗} Deutsches Fortschungszentrum für Künstliche Intelligenz Stuhlsatzenhausweg 3 D-6600 Saarbrücken 11 busemann@dfki.uni-sb.de

Recently the problem of choosing communicatively adequate lexemes has attracted much interest in the NL generation community. The problem is so fascinating because it includes a variety of subproblems that relate to all parts of the generation process. Up to now, many of the subproblems have been identified, and solutions have been suggested for each. However, steps towards an integrating approach to lexical choice are still missing.

Whether lexical choice must be exact in the sense that all and only the intended meaning is conveyed, depends on the respective communication situation. In a multimodal discourse, where language is supplemented by gestures or graphics, the linguistic device need not convey everything to the partner. In written discourse without shared contextual knowledge, exact verbalization seems much more in order. We will concentrate here on communication with NL as the only mode.

Lexical choice as a knowledge-based process

“Lexical choice” has become a general term for a variety of subproblems that involve the mapping from conceptual entities to lexical ones:

Definite reference: If a set of discourse referents is selected for verbalization, they must be described by concepts and, eventually, words that allow for an unambiguous identification of the referents. This includes the appropriate use of proforms.

Choice of open class words: Given a conceptual representation of the intended literal meaning, lexemes must be selected from a set of candidate lexemes that are communicatively appropriate in the given situation.

∗This research was supported by the German Federal Minister for Research and Technology under grant no. ITW 9002.

Collocations: There are different kinds of cooccurrence restrictions between lexemes. Some words cannot be used together with others (e.g. make a film but * make a photo), some tend to be used together with others (e.g. question and ask) and some yield a different meaning when used with certain others (idioms).

Connotations: Some words carry social connotations with them. German Putzfrau and Raumpflegerin both mean cleaning woman, but only the latter is used officially since the former carries a pejorative connotation with it. For related issues see (Hovy 1987).

Sublexical relations: Many concepts (words) have associative links to other concepts (words) that a speaker should be aware of or even can make use of. The concepts lead and heavy are related, since lead is heavy. The words flower and power are related due to their similar phonology (and probably also by their use in the late Sixties).

Work in lexical choice has concentrated so far on the choice of open class words and on collocational issues. Work on referring expressions was done for years, but it never seems to have been integrated with solutions to the two former problems.

What kinds of knowledge is lexical choice based on? The generation of faithful referring expressions relies on knowledge about the domain and the state of discourse, and on a dynamic model of the user’s goals and beliefs. The concepts expressed previously may influence subsequent choices.

The state of the discourse, especially the topic and focus constraints, influences the choice of open-class words. A moving event of a car turning from one street into another might be verbalized by either (1) or (2).

(1) The car turns into Hartungstreet.

(2) The car leaves Schluterstreet.

The choice of an open-class word can imply semantic and syntactic constraints for subsequent choices. The use of a verb such as to decide requires us to mention the agent (who decided) as well as information about the time of the action, which is not necessarily the case if we chose a corresponding noun such as decision.

(3) That John has decided to leave will be good for Mary.

(4) The decision to leave will be good for Mary.

Consequently, knowledge about different ways of syntactic realization must be available in order to ensure that the constraints imposed by the choice of various lexemes can still be satisfied by the grammar.

Likewise, collocational restrictions constrain follow-up choices. Often they amount to idiosyncratic language-specific properties of words making it difficult to generalize over them. A representation of collocations as lexical functions from words into words has been suggested (Mel’čuk, Polgu’ere 1987) that can be integrated in a lexical choice process (a suggestion within the framework of Systemic Functional Grammar is found in (Wanner, Bateman 1990)).

The representation of connotations and sublexical relations is being investigated in lexical semantics (for a recent overview see (Pustejovsky 1991)).

As it turns out, most of the knowledge sources generally required for NL generation are needed for the subtasks of lexical choice listed above. This stresses the fact that lexical choice should be recognized as a central process in NL generation that includes decisions about both what to say and how to say it. Viewing lexical choice as part of surface generation, as it has sometimes been suggested, is therefore inadequate.

Some results and observations

A brief sketch of some linguistic phenomena shows that lexical choice requires that the domain of discourse be modeled according to linguistic considerations.

Taxonomic relations. It should be uncontroversial that definite reference and pronominalization require taxonomic reasoning. (5) involves generalization (choosing a word whose meaning subsumes those of truck and Mercedes). As Reiter (this volume) points out, what is looked for is not the most specific item (e.g. motor-vehicle) but the most specific basic-level class (Rosch 1978). To solve this selection problem by classification, a domain model should include the property of concepts of being basic-level, which must, however, not be bequeathed.

(5) We’ll take the big truck and the Mercedes. Both vehicles are available.

(6) The big truck can’t be used. The engine doesn’t work.

Definite descriptions can also express various relations between two concepts, as in (6). By uttering the engine, the speaker assumes that the hearer can reconstruct (or establish) the has-part relation between truck and engine.

Syntactic constraints: Inalienable possessives in German allow definite descriptions without explicit prior introduction of the referent (7).^{1} While in (8) the owner of the ladder is verbalized as a genitive attribute, in (7) the owner of the foot is expressed as a dative NP.

(7) Hans trat Martha auf den Fuß. [Hans stepped Martha on the foot.]

(8) Hans trat auf Marthas Leiter. [Hans stepped on Martha’s ladder.]

Using the dative construction in the second case would lead to inacceptability. The genitive construction in the former case would lay the focus of attention on Martha, expressing that it was Martha’s foot which Hans stepped on (and not someone else’s). There are, however, no clear boundaries between what must go with a dative or with a genitive. With Auto [car], the dative construction can express the same meaning as the genitive (9).

1The brackets contain an interlinear translation.

(9) Hans trat Martha gegen das Auto. [Hans stepped Martha against the car.]

A domain model should exhibit the information which objects count as inalienable possessives.

Inferences. In a cooperative discourse, the speaker should avoid false communicative implicatures. (Reiter 1990) suggests an operationalization of some of the Gricean maxims and specifies conditions under which a speaker prefers a short utterance over a more verbose one and vice versa. For hearer-oriented lexical choice, two prerequisites are needed. First, the speaker’s communicative goal must be distinguished carefully from the speaker’s knowledge about the discourse topic. Second, a dynamic model of the hearer’s knowledge and of her concept-word relationships is important here (see Reiter, this volume for more detail). This user model is implemented by adding an “overlay” to the domain model (the latter is assumed to be shared by speaker and hearer). This overlay can be exchanged if a different hearer is addressed.

Events causing changes of state. The use of a pronoun is usually based on reference identity (or on a subsumption relation). A less well-defined relation is entailed in (10) where the antecedent referents of them undergo several changes of state according to the verbs grind and cook to a paste. This can be captured by keeping track of the changes and allowing reference to changed objects (Dale 1989). However, such an approach would have to explain the badness of (11).

(10) Grind the carots and potatos, cook them to a paste and put them into a prewarmed bowl.

(11) * Melt the ice and pour it into a glass.

The important interrelation between event representation and changes of state has only recently been thematized (Webber 1992).

Perspective. An event of a car changing its possessor in exchange for some money–assume the meaning being represented as in (12)–can be described by various lexically converse verbs, e.g. buy, sell, pay, receive etc. All participants in the event (buyer, seller, goods, money) can be verbalized, but not all are obligatory. If a participant is not verbalized, its existence in the event being described can be deduced. As (Fillmore 1977) notes, the verbs bring different participants into perspective. Verb choice must thus take into consideration which discourse referents are in perspective. This is determined by dynamic context knowledge. (Jacobs 1987) demonstrates how lexical and conceptual knowledge represented in one and the same formalism (called Ace) can be interrelated for verb choice.

(12) \begin{matrix}CommercialExchange(goods: car123, buyer: Peter, \\ seller: John, money: $800) \end{matrix}

The domain model must hence be capable of expressing entities in perspective as well as different views of some event (Jacobs 1987). A view may relate concepts to words. For instance, if the seller and the goods are in perspective, it is most appropriate to describe (12) by the verb sell. The informal notion of a view addresses the connection between the domain model and the lexicon, which may turn out to be complicated if different formalisms are used for the two knowledge sources.

Dimensional expressions. On the basis of the defining features of objects alone it is not possible to describe an object’s maximal axis. Take e.g. a pole: whether (13) or (14) is appropriate depends on the pole’s position given by the context; only if the pole is upright, (13) is possible.

(13) The pole is 10 m high.

(14) The pole is 10 m long.

(Lang, Carstensen, Simmons 1990) suggest a two-staged propositional semantic representation that relates language-independant conceptual entities to lexical entities. Among other things, conceptual features of spatial objects (pole has a maximal axis) are combined with lexical constraints of dimensional adjectives (high requires, in contrast to long, a vertical orientation of an object’s maximal axis).

The selection of dimensional adjectives also depends on the speaker’s spatial location. (Lang, Carstensen, Simmons 1990) claim that their theory purports to any dimensional expression, thus including prepositions, nouns, and verbs as well.

For this approach to be used in a lexical choice system, the model of spatial knowledge must be geared towards distinguishing the two levels carefully.

Prerequisites for an integrating framework

The above-mentioned phenomena are manifold. A unifying framework for their treatment has not been developed yet. The only step in this direction has been undertaken by (Nirenburg et al. 1988) who suggest a blackboard architecture to deal with definite descriptions, ellipsis, pronominalization, and collocational restrictions. All other approaches are designed to cope with a small subset of the above lexical-choice phenomena. The need for a careful distinction between a conceptual and a lexical level is obvious in every case. Hence an attempt to build a unifying framework for lexical choice, which can capture a broader spectrum of phenomena should be based on explicit relations between concepts and words (cf. Novak, this volume).

Moreover, a choice of an open-class word should rely on information about the hearer’s knowledge, on her capabilities to relate words to concepts and on the current situational and linguistic context. Finally, the syntactic and stylistic consequences emerging from a decision should be considered.

An integrating framework must include a powerful knowledge representation formalism that allows us to model, besides subsumption and relations (roles) the different special constructs, such as the basic-level property, inalienable possessives, perspective, changes of state, views, etc.

Three ways of relating conceptual to lexical entities can be pursued. The most common one keeps the lexicon separate from conceptual knowledge. This acknowledges the different nature and use of concepts and words. However, this would involve the definition of a complex interface (remember that a choice may fail due to constraints in either part).

Alternatively, the lexicon could be represented using the same knowledge representation formalism as for the domain model. This has been realized e.g. in Ace (Jacobs 1987). However, the content and structure of a lexicon depends on the chosen grammar formalism. This may require different representational means. For instance, a constraintbased approach such as HPSG (Pollard, Sag 1987) uses mechanisms for inheritance that extend those of KL-ONE.

Third, work in lexical semantics could be taken up. (Pustejovsky 1991) suggests to incorporate taxonomic distinctions into the lexicon. To my knowledge this path has not been followed yet for lexical choice. It requires, in practice, the reconstruction of knowledge representation techniques within the lexicon formalism.

Which approach is best taken depends on the yet open question of how the information needed for lexical choice should best be divided up between the lexicon and the domain model.

Literatur

Dale, R. (1989): Cooking up referring expressions. In: Proceedings of the 27th Annual Meeting of the Association for Computational Linguistics, Vancouver, BC., Canada

Fillmore, C. J. (1977): The case for case reopened. In: Cole, P., Sadock, J. M. (eds.): Syntax and Semantics Vol. 8. Grammatical Relations. Academic Press, New York

Hovy, E. H. (1987): Generating Natural Language Under Pragmatic Constraints. PhD thesis, Department of Computer Science, Yale University

Jacobs, P. (1987): KING: a knowledge-intensive natural language generator. In: Kempen, G. (ed.): Natural Language Generation. Martinus Nijhoff Publishers, Dordrecht, Boston, Lancaster, pp. 219–230. NATO ASI Series E Nr. 135

Lang, E., Carstensen, K.-U., Simmons, G. (1990): Modelling Spatial Knowledge on a Linguistic Basis. IWBS Report 147, IBM Deutschland, Institut f”ur wissensbasierte Systeme, Postfach 80 08 80, 7000 Stuttgart 80

Mel’čuk, I., Polgu’ere, A. (1987): A Formal Lexicon in the Meaning-Text-Theory (or How to Do Lexica with Words). Computational Linguistics 13, 261–275

Nirenburg, S., McCardell, R., Nyberg, E., Werner, P., Huffman, S., Kenschaft, E., Nirenburg, I. (1988): DIOGENES-88. Report CMU-CMT-88-107, Carnegie Mellon University, Center for Machine Translation

Pollard, C. J., Sag, I. A. (1987): Information-Based Syntax and Semantics. Vol. 1. Univ. of Chicago Press, Chicago. Center for the Studies of Language and Information, CSLI Lecture Note No. 13

Pustejovsky, J. (1991): The Generative Lexicon. Computational Linguistics 17(4), 409–441

Reiter, E. B. (1990): A New Model for Lexical Choice for Open-Class Words. In: Proceedings of the Fifth International Natural Language Generation Workshop, Dawson, Pa., pp. 23–30

Rosch, E. (1978): Principles of Categorization. In: Rosch, E., Lloyd, B. (eds.): Cognition and Categorization. Lawrence Erlbaum, Hillsdale, pp. 27–48

Wanner, L., Bateman, J. (1990): A Collocational Based Approach to Salience-Sensitive Lexical Selection. In: Proceedings of the Fifth International Natural Language Generation Workshop, Dawson, Pa., pp. 31–38

Webber, B. L. (1992): Accommodating Context Change. In: Proc. Conf. of the 30th Annual Meeting of the ACL, Newark